OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 2.8x = 5.1 A. 1.7 B. 1.8 C. 1.9 D. 18.2 I got B.

OpenStudy (anonymous):

2.8x = 5.1 x = 5.1/2.8 x = 1.821 rounded to nearest tenth = 1.8 you are correct.....answer is B

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.15(x – 8.86) = 21.83

OpenStudy (anonymous):

i got 12.7

OpenStudy (anonymous):

A. 12.8 B. 12.7 C. 12.6 D. 11.7

OpenStudy (anonymous):

you still there

OpenStudy (anonymous):

6(x - 9) = 22 6x - 54 = 22 6x = 22 + 54 6x = 76 x = 76/6 = 12.666 rounds to 12.7 Not sure if it is 12.66 rounded to 12.7 or if they don't want the answer rounded, it is 12.6 ??

OpenStudy (anonymous):

Which estimate is most reasonable? Francesca purchases a box of cereal for $2.98, hamburger for $4.12, tomatoes for $3.07, and milk for $1.96. She has $17 with her. About how much change should she expect to receive? A. $7 B. $6 C. $5 D. $4

OpenStudy (anonymous):

I would probably round it to 12.7......so I agree with you again

OpenStudy (anonymous):

3 + 4 + 3 + 2 = 12 17 - 12 = 5 --> answer C

OpenStudy (anonymous):

I think it is $4 but calculated it is 4.87 and it would be rouded to 4.87

OpenStudy (anonymous):

round your prices...... 2.98 = 3 4.12 = 4 3.07 = 3 1.96 = 2 added together = 12 subtract from 17 = 5

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

need any more help ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I can try to help

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 3.6x = 5.5 A. 15.3 B. 1.6 C. 1.5 D. 1.4

OpenStudy (anonymous):

i got 1.5

OpenStudy (anonymous):

rounded to the nearest tenth

OpenStudy (anonymous):

3.6x = 5.5 x = 5.5/3.6 x = 1.52 rounds to 1.5 your correct

OpenStudy (anonymous):

i need your help with 4 more questions

OpenStudy (anonymous):

lets see what you got ?

OpenStudy (anonymous):

Solve the equation. Round the answer to the nearest tenth. 1.18x + 4.34 = 8.67 A. 11.0 B. 3.8 C. 3.7 D. 3.6

OpenStudy (anonymous):

i got 3.7

OpenStudy (anonymous):

kelliegirl

OpenStudy (anonymous):

your correct, it is 3.7...sorry about that, my computer clicked off. I am back now

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.2x + 1.2 = 8.9 A. 1.3 B. 1.4 C. 1.6 D. 1.7

OpenStudy (anonymous):

6x + 1 = 9 6x = 9 - 1 6x = 8 x = 8/6 = 1.33 rounds to 1.3

OpenStudy (anonymous):

i got that too

OpenStudy (anonymous):

great minds think alike :)

OpenStudy (anonymous):

ha

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 5.2x + 3.2 = 9.8 A. 1.5 B. 1.1 C. 1.2 D. 1.4

OpenStudy (anonymous):

5x + 3 = 10 5x = 10 - 3 5x = 7 x = 7/5 = 1.4

OpenStudy (anonymous):

wow i got that answer

OpenStudy (anonymous):

Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 6.2x + 1.2 = 8.9 A. 1.3 B. 1.4 C. 1.6 D. 1.7

OpenStudy (anonymous):

6x + 1 = 9 6x = 9 - 1 6x = 8 x = 8/6 = 1.33 rounds to 1.3 wait a minute.....we already did this one

OpenStudy (anonymous):

sorry i did the wrong one here Which is the best estimate of x based on rounding the constants and coefficients in the equation to the nearest integer? 3.09(x – 6.91) = 24.95 A. 15.2 B. 10.3 C. 15.4 D. 15.3

OpenStudy (anonymous):

3(x - 7) = 25 3x - 21 = 25 3x = 25 + 21 3x = 46 x = 46/3 = 15.33 rounds to 15.3

OpenStudy (anonymous):

got that

OpenStudy (anonymous):

you know this stuff pretty good :)

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

Which estimate is most reasonable? Michaela purchases a box of cereal for $3.95, chicken for $5.99, orange juice for $2.10, and a loaf of bread for $1.05. She has $15 with her. About how much change should she expect to receive? A. $1 B. $3 C. $4 D. $2

OpenStudy (anonymous):

i got $1.91

OpenStudy (anonymous):

4 + 6 + 2 + 1 = 13 15 - 13 = 2 <-- answer D

OpenStudy (anonymous):

round your prices, it makes it easier

OpenStudy (anonymous):

thank you so very much that was all

OpenStudy (anonymous):

glad I could help :)

OpenStudy (anonymous):

@kelliegirl33 i am a fan of you know on here

OpenStudy (anonymous):

thats fine with me :)

OpenStudy (anonymous):

ok bye

OpenStudy (anonymous):

bye.....have a good day :)